OpenStudy (anonymous):

Find S15 for the geometric series 72 + 12 + 2 + +…

OpenStudy (anonymous):

This is what we just did, but notice what is happening to the terms. What is r here?

OpenStudy (anonymous):

72-12=60 12-2=10

OpenStudy (anonymous):

Geometric means we multiply or divide to go from one term to the next so.....

OpenStudy (anonymous):

R=6

OpenStudy (anonymous):

1/6, we are dividing to go from one term to the next. See?

OpenStudy (anonymous):

So the equation is 72(1-6^15)/(1-6)

OpenStudy (anonymous):

\[72 (\frac{ 1 - (\frac{ 1}{ 6 })^{15} }{ 1-\frac{ 1 }{ 6 } })\]

OpenStudy (anonymous):

Careful calculating. Do it one little piece at a time. What calculator are you using?

OpenStudy (anonymous):

A Computer Calculator

OpenStudy (anonymous):

You answer will be a "nice" number. Not a ton of decimal places.